OpenStudy (anonymous):

What is the value of x? And what is the measure of angle 1?

OpenStudy (anonymous):

Heres the pic, I really need help -.-

OpenStudy (anonymous):

x + 30 + 4x = 90 5x + 30 = 90 5X = 60 X = 12 Plug x in to each angle to find measure, then all four angles add to equal 180

OpenStudy (anonymous):

First part answer choices 10o 25o 30o 120o What is the measure of 1? answer choices 45o 60o 120o 125o

OpenStudy (anonymous):

x+30+4x=180 ------- alternating angle & angle on straight line 5x+30=180 5x=150 x=150/5 x=30

OpenStudy (anonymous):

I understand now :D